A 270 kg satellite is orbiting on a circular orbit 6180 km above the Earth's surface. Determine the speed of the satellite. (The mass of the Earth is 5.97×1024 kg, and the radius of the Earth is 6370 km.)
Answer #1

here,

the mass of satellite , m = 270 kg

altitude of orbit , A = 6180 km = 6.18 * 10^6 m

mass of earth , M = 5.97 * 10^24 kg

radius of earth , R = 6370 km = 6.37 * 10^6 m

the speed of satellite , v = sqrt(G * M /(R + A))

v = sqrt(6.67 * 10^-11 * 5.97 * 10^24 /(6.18 * 10^6 + 6.37 * 10^6))

v = 5.63 * 10^3 m/s

the speed of satellite is 5.63 * 10^3 m/s